Suppose you have a 12.0 V car battery that can move 45 525 C of charge, and a 12.0 V motorcycle battery that can move 3250 C of charge. How much energy does each deliver?





1
Expert's answer
2022-05-25T10:27:31-0400
"W_1=q_1V_1=45525*12.0=5.46*10^5\:\rm J"

"W_2=q_2V_2=3250*12.0=3.90*10^4\:\rm J"